# Synthetic Homes: An Accessible Multimodal Pipeline for Producing Residential Building Data with Generative AI

### v2

@@ Abstract
Computational models have emerged as powerful tools for energy modeling research, touting scalability and quantitative results. However, these models require a plethora of data, some of which can be inaccessible, expensive, or can raise privacy concerns. We introduce a modular multimodal framework to produce this data from publicly accessible images and residential information using generative Artificial Intelligence (AI). Additionally, we provide a pipeline demonstrating this framework and we evaluate its generative AIcomponents. Our experiments show that our framework’s use of AIavoids common issues with generative models and produces realistic multimodal data. By reducing dependence on costly or restricted data sources, we pave a path towards more accessible research in Machine Learning (ML)and other data-driven disciplines.
keywords:
Building Energy Models (BEM) , Urban Building Energy Modeling (UBEM) , Energy Policy , Synthetic Data , Generative AI
journal: Journal of Building Engineering
[…]
[label1]organization=Department of Computer Science, Lafayette College, addressline=730 High Street, city=Easton, state=Pennsylvania, postcode=18042, country=United States of America
[…]
[label2]organization=Departments of Geosciences & Computer Science, Georgia State University, addressline=38 Peachtree Center Ave. SE, Suite 730, city=Atlanta, state=Georgia, postcode=30303, country=United States of America
[…]
[label3]organization=Center for Disaster Informatics & Computational Epidemiology, Georgia State University, addressline=38 Peachtree Center Ave. SE, Suite 730, city=Atlanta, state=Georgia, postcode=30303, country=United States of America
{graphicalabstract}[Uncaptioned image]
{highlights}
Novel modular framework to generate synthetic data for use in urban energy research.
Uses textual and visual analysis for generating data with LLMs for direct use in BEM.
Data generated is validated to reduce biases that result from the use of LLMs.
Includes statistical evidence that shows generated data is realistic.
Produces data while spending less than US \$1 per home for GPT usage and electricity.
@@ 1 Introduction
Consumption of electricity in the United States will increase by $1.7\%$ between 2020 and 2026 [1] and seasonal electricity use per person has nearly doubled from 1973 to 2025 [2]. Rising energy consumption exerts increasing pressure on infrastructure capacity, operational costs, and greenhouse gas emissions, resulting in rising costs for governments, energy producers, and energy consumers. Accurate modeling and policy tools are therefore essential for effective planning, aiding by preventing infrastructure strain, optimizing energy costs, minimizing or mitigating climate impacts, assessing policy effectiveness, and guiding decision-making in urban centers [3, 4, 5, 6].
Computational approaches to monitor energy usage and model urban centers, buildings, and other structures require large amounts of data, such as type and quality of materials, floor plans, and weather patterns to produce scalable and quantitative results [7]. However, several issues arise when attempting to collect such information, including high prices, limited availability, and privacy concerns. These issues create barriers to accessing data, hindering the development of accurate models and the scaling of models to larger populations or geographic areas. The inaccessibility of this data has led researchers to generate synthetic data through a variety of methods. We propose a generative AIbased approach to generate low-cost multimodal synthetic data consisting of simulations and home inspection notes that can be used to develop, validate, and improve existing computational models for energy demand modeling. The proposed method enables the generation of large volumes of data for use in any geographic region where basic descriptive data on housing structures is publicly available.
@@ 2 Related Works
Current data generation efforts span a wide range of approaches, from optimizing building construction processes to the creation of digital twins. For example, Wan et al. [8] built a generative model that improves the efficiency of energy consumption by determining the optimal layout of floor plans. Elnabawi et al. [9] enhanced their building energy models using highly detailed microclimate models that considered features such as shadow generation and seasons to generate weather data. Lee et al. [10] train a neural network that predicts building energy consumption, aiming to create a more robust model with limited real-world measurements. The network was trained with data synthesized using, among others, electricity, temperature, and humidity, encoded into receiver operating characteristic (ROC) plots which visually encode the separation between classes and are commonly used to assess model discrimination [11, 12].
@@
Other researchers focus on creating and applying digital twins: virtual, dynamic, and data-driven digital replicas of real-world urban environments. Stinner et al. [13] generate digital twins of building energy systems from piping and instrumentation diagrams using convolutional networks and other algorithms. Agostinelli et al. [14] create digital twins of energy efficient buildings in residential areas from Building Information Modeling (BIM), Internet of Things (IoT), Geographical Information Systems (GIS), and AIcomponents, with machine learning serving as the central module of the model. Francisco et al. [15] take a different approach by generating digital twins of daily, time-segmented energy benchmarks solely from smart meter data streams. Belik and Rubanenko [16] present a hybrid digital twin framework for renewable energy sources (RES) that integrates real-time data, 3D modeling, and sensitivity analysis to improve forecasting and operational efficiency, particularly in grid systems with high RES penetration.
@@
While prior work has focused on traditional simulation tools or sensor data to construct digital twins, Xu et al. [17] encourage the use of generative AImodels, such members of the GPTfamily, because they present new opportunities in urban data generation. For instance, Dodge et al. [18] employ convolutional neural networks (CNNs) to parse floor plan images, generating structured representations using wall segmentation, object recognition, and optical character recognition. Relatedly, Zhang et al. [19] implement a pipeline that generates EnergyPlusIDF files from initial building geometry and textual building descriptions. Moreover, Xiao and Xu [20] propose a multi-agent Large Language Model (LLM)framework that processes unstructured building data and sensor logs to support automated energy optimization in smart grid contexts. Additionally, in closely related fields, generative AIhas been successfully employed to perform a variety of tasks, such as predictive analysis of traffic congestion [21], real-time decision-making in urban air mobility [22], and stochastic solar irradiance forecasting for building facades [23].
@@
Liu et al. [24] present a wide-ranging review on the integration of generative AIinto building energy workflows, highlighting both opportunities and challenges. These include concerns such as dataset preparation, hallucinations, and a lack of domain-specific expertise. However, the impact of such challenges varies depending on the pipeline design. Because our framework combines pretrained LLMswith simulation tools for quantitative evaluation, we avoid the need for fine-tuning or physics-aware generation, thereby reducing the impact of these concerns. We evaluate our framework’s LLMcomponents to reduce hallucinations. Additionally, Liu emphasizes the importance of multimodal integration—where multiple data modalities (such as image, text, and tabular data) are used by a process or model. Producing high-quality structured urban data suitable for tasks such as modeling and simulation requires combining multiple complementary processes into a single workflow [25, 26, 27]. We adopt these principles in the design of our general framework.
@@ 3 Methodology
    An image processor powered by LLaVA [28];
@@
    A Geographic JSON (GeoJSON)and inspection note generator using GPT [29]; and
@@
    An EnergyPlus [^1] EnergyPlus™ is a trademark of the United States Department of Energy. [30] simulation
@@
An instance of this pipeline was implemented and executed [^2] Our source code is available at: https://github.com/Lafayette-EshbaughSilveyra-Group/synthetic-homes. on a g3.medium flavor JetStream 2 virtual machine running Ubuntu 24.04.2 (Noble). Specifically, the system featured one NUMA node with 8 AMD EPYC-Milan processors, 30 GB of RAM, and an NVIDIA A100 SXM4 40 GB GPU (of which 10 GB was allocated). Additionally, it was fully virtualized (KVM and AMD-V). We paid \$0.36 (USD) for OpenAI requests to produce data for 258 homes—about \$0.0014 per home. [^3] Assuming a conservative upper bound of 10 CPU minutes per home at 100 watts and electricity prices at \$0.15 per kilowatt-hour, local electricity costs are approximately \$0.0025 per synthetic home, giving a worst case total of \$0.0039 per synthetic home. To run the pipeline, we used the PyTorch [31] machine learning framework, and the Hugging Face Transformers library [32] to load the models we used. In the subsections that follow, we describe each pipeline component in detail.
@@ 3.1 Scraping the Data
To ensure the synthetic data we generated reflects real-world homes, we collected data from publicly available county datasets. For our pipeline instance, we gathered data from Northampton County, Pennsylvania, where two of the authors work and reside. To collect this data, we implemented a web scraper using Selenium WebDriver [33] and ChromeDriver [34]. For each residential property on each street in the list of streets to process, the scraper extracts key attributes, such as number of rooms, total floor area, and other metadata listed in Table 1. This data is saved to a JavaScript Object Notation (JSON)file for each home. Additionally, the scraper downloads two images: a street view photograph and a floor plan, both of which are made available on the County’s public-facing platform. An example of each is depicted in Figure 2.
Figure 2: An example street view photograph and floor plan from the Northampton County database.
@@ 3.2 Processing Images
After scraping the street view photographs and floor plans from the county portal, they are analyzed using LLaVA [28]. This step processes both files to translate them into a textual description. For example, the floor plans can be used to define the geometry of the home, and the image to determine the quality and number of windows. The use of images allows our pipeline to infer important aspects of the homes that will be used to generate the GeoJSON, which is ultimately used to generate the Input Definition File (IDF)file required by EnergyPlus.
@@ 3.2.1 Evaluating Visual Focus with Occlusion Testing
Occlusionmeasures the necessity of a part of an image: if removing a feature causes the prediction to change, that feature was necessary for the model’s decision [35, 36]. When we remove image portions, we replace them with a neutral value, such as black. This creates a masked version of the image. The degree of change measured between the model’s output for this masked image and its baseline response reflects the importance of the masked cell. High sensitivity to the removal of a feature suggests the model relied heavily on that region to produce its output, indicating that the region was internally necessary to the model’s decision process.
During the development of our pipeline, we identified two candidate LLMsto employ for image processing. To determine the best model for this task, we compared the focus of the GPTand LLaVALLMswhen analyzing images using occlusion tests. We provided the models with twenty different images of homes: ten with roofs showing no signs of damage, and ten with roofs with significant damage. Each model was instructed to evaluate and describe the state of the roof and justify its evaluation. We first took a baseline where no part of the image was masked. Then, the image was divided into one hundred cells, a number we selected to balance effectiveness, efficiency, and cost. Each cell of the image was masked, and the model was executed on the original prompt and this new image. Textual responses were embedded into vectors using the all-MiniLM-L6-v2 SentenceTransformers model [37], a lightweight and widely used model sufficient for obtaining sentence-level numerical representations. Then, to measure the semantic difference between each execution of the experiment, we took the cosine distance of each embedded response from the baseline. Additionally, heatmaps of these differences were assembled using matplotlib [38].
@@ 3.3 Generating GeoJSON
We use GPTfor structured data generation given its strong ability to produce well-formed JSONoutputs from descriptive prompts. Multiple studies have shown that LLMsperform well on generation tasks under strict schema constraints, even with no previous prompting or fine tuning. For example, StructuredRAG’s benchmark suite showed $82\%$ format compliance under JSONprompt conditions [39]. As we built the pipeline, we reviewed GPT’s responses and it seldom failed to adhere to the JSONstructure. In the rare case that it doesn’t, our program catches the JSONerror and automatically reprompts GPTfor that example.
@@ 3.4 Running EnergyPlus Simulations
The templating process is facilitated with the eppy package [40], which makes programmatic IDFgeneration straightforward. Then, we use ExpandObjects to prepare the file for simulation. ExpandObjects is bundled with EnergyPlus, and its primary use case is to take an IDFand expand template objects—such as the HVACTemplate we use—into full systems [41]. Finally, we run the expanded IDFfile through EnergyPlus, producing final simulation results.
@@ 4 Validating Realism of Pipeline Generated Data
Following prior work on synthetic data evaluation [42, 43], we adopt the definition of realism as the overlap of synthetic values with the $10\%$$90\%$ range of a trusted reference on the home characteristics being generated. To measure the realism of our generated data, we compare it with ResStock [44, 45], a dataset that contains realistic samples of homes in different regions of the United States as IDFfiles. The analysis consists of performing statistical analyses comparing the distribution of values in our synthetic homes to those in ResStock. We only use ResStock data from climate zone 4A, the region in which the synthetic homes are collected in our example, and we only compare the four pre-selected variables—Wall R-Value, Roof R-Value, Cooling COP, Heating COP—that are not static among all synthetic homes. The number of synthetic homes is $258$ while the number of climate zone 4A ResStock samples is $2000$.
@@
The results above and the tests of the generative components of our framework demonstrate a level of confidence above previous work such as Zhang et al. [19] and Xiao and Xu [20]. These works and others either accept the output of the LLMsor use procedural evaluation techniques.
Occlusion offers stronger empirical support that the image processing component of our pipeline—which is LLM-based—performs as expected. We additionally demonstrated statistically that properties of our generated homes are realistically distributed according to ResStock. Our results directly address and mitigate concerns about hallucination and unreliability raised in prior work [24]. It is fair to be skeptical of LLMsand other MLmodels—but with our results, we clearly demonstrate our pipeline is not as susceptible to some of the issues Liu et al. [24] raise.

### v3

@@ Abstract
Computational models have emerged as powerful tools for multi-scale energy modeling research at the building level as well as urban scale. However, these models require a plethora of data on building parameters, some of which can be inaccessible, expensive, or can raise privacy concerns. We introduce a modular multimodal framework to synthetically produce this data from publicly accessible images and residential information using generative Artificial Intelligence (AI). Additionally, we provide a modeling pipeline demonstrating this framework and we evaluate its generative AIcomponents for realism. Our experiments show that our framework’s use of AIavoids common issues with generative models and produces realistic multimodal data at the building scale. Resulting datasets can be used for assessing influence of energy efficiency upgrades at the building scale, as well as to simulate larger patterns of energy consumption across regions. This work will support research in building and energy simulation by reducing dependence on costly or restricted data sources, and pave a path towards more accessible research in Machine Learning (ML)and other data-driven disciplines.
Keywords: Building Energy Models (BEM), Urban Building Energy Modeling (UBEM), Energy Policy, Synthetic Data, Generative AI
@@ 1 Introduction
Consumption of electricity in the United States will increase by $1.7\%$ between 2020 and 2026 (U.S. Energy Information Administration, 2025) and seasonal electricity use per person has nearly doubled from 1973 to 2025 (U.S. Environmental Protection Agency, 2025). Rising energy consumption, due to inefficient building mechanics, exerts increasing pressure on infrastructure capacity, operational costs, and greenhouse gas emissions, resulting in rising costs for governments, energy producers, and energy consumers. Accurate modeling and policy tools are therefore essential for effective planning, aiding by preventing infrastructure strain, optimizing energy costs, minimizing or mitigating climate impacts, assessing policy effectiveness, and guiding decision-making in urban centers (Bompard et al., 2024; Perera et al., 2021; Zeng et al., 2011; Hajri et al., 2025).
Computational approaches—such as the use of Building Energy Models (BEMs)to monitor energy usage and model urban centers, buildings, and other structures—require large amounts of data, such as type and quality of materials, floor plans, and weather patterns to produce scalable and quantitative results (U.S. Department of Energy, ). However, several issues arise when attempting to collect such information, including high prices, limited availability, and privacy concerns. These issues create barriers to accessing data, hindering the development of accurate models and the scaling of models to larger populations or geographic areas. The inaccessibility of this data has led researchers to generate synthetic data through a variety of methods. We propose a generative AIbased approach to generate low-cost multimodal synthetic data consisting of simulations and home inspection notes that can be used to develop, validate, and improve existing computational models for energy demand modeling. The proposed method enables the generation of large volumes of data for use in any geographic region where basic descriptive data on housing structures is publicly available.
@@ 2 Related Works
Current data generation efforts span a wide range of approaches, from optimizing building construction processes to the creation of digital twins. For example, Wan et al. (2022) built a generative model that improves the efficiency of energy consumption by determining the optimal layout of floor plans. Elnabawi and Hamza (2022) enhanced their building energy models using highly detailed microclimate models that considered features such as shadow generation and seasons to generate weather data. Lee et al. (2019) train a neural network that predicts building energy consumption, aiming to create a more robust model with limited real-world measurements. The network was trained with data synthesized using, among others, electricity, temperature, and humidity, encoded into receiver operating characteristic (ROC) plots which visually encode the separation between classes and are commonly used to assess model discrimination (Zweig and Campbell, 1993; Janssens and Martens, 2020).
@@
Other researchers focus on creating and applying digital twins: virtual, dynamic, and data-driven digital replicas of real-world urban environments. Stinner et al. (2021) generate digital twins of building energy systems from piping and instrumentation diagrams using convolutional networks and other algorithms. Agostinelli et al. (2021) create digital twins of energy efficient buildings in residential areas from Building Information Modeling (BIM), Internet of Things (IoT), Geographical Information Systems (GIS), and AIcomponents, with machine learning serving as the central module of the model. Francisco et al. (2020) take a different approach by generating digital twins of daily, time-segmented energy benchmarks solely from smart meter data streams. Belik and Rubanenko (2023) present a hybrid digital twin framework for renewable energy sources (RES) that integrates real-time data, 3D modeling, and sensitivity analysis to improve forecasting and operational efficiency, particularly in grid systems with high RES penetration.
@@
While prior work has focused on traditional simulation tools or sensor data to construct digital twins, Xu et al. (2024) encourage the use of generative AImodels, such members of the GPTfamily, because they present new opportunities in urban data generation. For instance, Dodge et al. (2017) employ convolutional neural networks (CNNs) to parse floor plan images, generating structured representations using wall segmentation, object recognition, and optical character recognition. Relatedly, Zhang et al. (2025a) implement a pipeline that generates EnergyPlusIDF files from initial building geometry and textual building descriptions. Moreover, Xiao and Xu (2024) propose a multi-agent Large Language Model (LLM)framework that processes unstructured building data and sensor logs to support automated energy optimization in smart grid contexts. Additionally, in closely related fields, generative AIhas been successfully employed to perform a variety of tasks, such as predictive analysis of traffic congestion (Lin et al., 2025), real-time decision-making in urban air mobility (Sha et al., 2024), and stochastic solar irradiance forecasting for building facades (Zhang et al., 2022).
@@
Liu et al. (2025) present a wide-ranging review on the integration of generative AIinto building energy workflows, highlighting both opportunities and challenges. These include concerns such as dataset preparation, hallucinations, and a lack of domain-specific expertise. However, the impact of such challenges varies depending on the pipeline design. Because our framework combines pretrained LLMswith simulation tools for quantitative evaluation, we avoid the need for fine-tuning or physics-aware generation, thereby reducing the impact of these concerns. We evaluate our framework’s LLMcomponents to reduce hallucinations. Additionally, Liu emphasizes the importance of multimodal integration—where multiple data modalities (such as image, text, and tabular data) are used by a process or model. Producing high-quality structured urban data suitable for tasks such as modeling and simulation requires combining multiple complementary processes into a single workflow (Rehmann et al., 2025; Guo et al., 2023; Bishop et al., 2024). We adopt these principles in the design of our general framework.
@@ 3 Methodology
    An image processor powered by LLaVA (Liu et al., 2023);
@@
    A Geographic JSON (GeoJSON)and inspection note generator using GPT (OpenAI, 2025); and
@@
    An EnergyPlus (Crawley et al., 2001) simulation
@@
An instance of this pipeline was implemented and executed [^1] Our source code is available at: https://github.com/Lafayette-EshbaughSilveyra-Group/synthetic-homes. on a g3.medium flavor JetStream 2 virtual machine running Ubuntu 24.04.2 (Noble). Specifically, the system featured one NUMA node with 8 AMD EPYC-Milan processors, 30 GB of RAM, and an NVIDIA A100 SXM4 40 GB GPU (of which 10 GB was allocated). Additionally, it was fully virtualized (KVM and AMD-V). We paid \$0.36 (USD) for OpenAI requests to produce data for 258 homes—about \$0.0014 per home. [^2] Assuming a conservative upper bound of 10 CPU minutes per home at 100 watts and electricity prices at \$0.15 per kilowatt-hour, local electricity costs are approximately \$0.0025 per synthetic home, giving a worst case total of \$0.0039 per synthetic home. To run the pipeline, we used the PyTorch (Ansel et al., 2024) machine learning framework, and the Hugging Face Transformers library (Wolf et al., 2020) to load the models we used. In the subsections that follow, we describe each pipeline component in detail.
@@ 3.1 Scraping the Data
To ensure the synthetic data we generated reflects real-world homes, we collected data from publicly available county datasets. For our pipeline instance, we gathered data from [location withheld for double-blind review]. To collect this data, we implemented a web scraper using Selenium WebDriver (Selenium, 2024) and ChromeDriver (Google, 2023). For each residential property on each street in the list of streets to process, the scraper extracts key attributes, such as number of rooms, total floor area, and other metadata listed in Table 1. This data is saved to a JavaScript Object Notation (JSON)file for each home. Additionally, the scraper downloads two images: a street view photograph and a floor plan, both of which are made available on the county’s public-facing platform. An example of each is depicted in Figure 2.
Figure 2: An example street view photograph and floor plan from the county database.
@@ 3.2 Processing Images
After scraping the street view photographs and floor plans from the county portal, they are analyzed using LLaVA (Liu et al., 2023). This step processes both files to translate them into a textual description. For example, the floor plans can be used to define the geometry of the home, and the image to determine the quality and number of windows. The use of images allows our pipeline to infer important aspects of the homes that will be used to generate the GeoJSON, which is ultimately used to generate the Input Definition File (IDF)file required by EnergyPlus.
@@ 3.2.1 Evaluating Visual Focus with Occlusion Testing
Occlusionmeasures the necessity of a part of an image: if removing a feature causes the prediction to change, that feature was necessary for the model’s decision (Hooker et al., 2019; Balkir et al., 2022). When we remove image portions, we replace them with a neutral value, such as black. This creates a masked version of the image. The degree of change measured between the model’s output for this masked image and its baseline response reflects the importance of the masked cell. High sensitivity to the removal of a feature suggests the model relied heavily on that region to produce its output, indicating that the region was internally necessary to the model’s decision process.
During the development of our pipeline, we identified two candidate LLMsto employ for image processing. To determine the best model for this task, we compared the focus of the GPTand LLaVALLMswhen analyzing images using occlusion tests. We provided the models with twenty different images of homes: ten with roofs showing no signs of damage, and ten with roofs with significant damage. Each model was instructed to evaluate and describe the state of the roof and justify its evaluation. We first took a baseline where no part of the image was masked. Then, the image was divided into one hundred cells, a number we selected to balance effectiveness, efficiency, and cost. Each cell of the image was masked, and the model was executed on the original prompt and this new image. Textual responses were embedded into vectors using the all-MiniLM-L6-v2 SentenceTransformers model (Reimers and Gurevych, 2019), a lightweight and widely used model sufficient for obtaining sentence-level numerical representations. Then, to measure the semantic difference between each execution of the experiment, we took the cosine distance of each embedded response from the baseline. Additionally, heatmaps of these differences were assembled using matplotlib (The Matplotlib Development Team, 2025).
@@ 3.3 Generating GeoJSON
We use GPTfor structured data generation given its strong ability to produce well-formed JSONoutputs from descriptive prompts. Multiple studies have shown that LLMsperform well on generation tasks under strict schema constraints, even with no previous prompting or fine tuning. For example, StructuredRAG’s benchmark suite showed $82\%$ format compliance under JSONprompt conditions (Shorten et al., 2024). As we built the pipeline, we reviewed GPT’s responses and it seldom failed to adhere to the JSONstructure. In the rare case that it doesn’t, our program catches the JSONerror and automatically reprompts GPTfor that example.
@@ 3.4 Running EnergyPlus Simulations
The templating process is facilitated with the eppy package (santoshphilip, 2024), which makes programmatic IDFgeneration straightforward. Then, we use ExpandObjects to prepare the file for simulation. ExpandObjects is bundled with EnergyPlus, and its primary use case is to take an IDFand expand template objects—such as the HVACTemplate we use—into full systems (U.S. Department of Energy, 2023). Finally, we run the expanded IDFfile through EnergyPlus, producing final simulation results.
@@ 4 Validating Realism of Pipeline Generated Data
Following prior work on synthetic data evaluation (Goncalves et al., 2020; Patki et al., 2016), we adopt the definition of realism as the overlap of synthetic values with the $10\%$$90\%$ range of a trusted reference on the home characteristics being generated. To measure the realism of our generated data, we compare it with ResStock (Present et al., 2024; Reyna et al., 2025), a dataset that contains realistic samples of homes in different regions of the United States as IDFfiles. The analysis consists of performing statistical analyses comparing the distribution of values in our synthetic homes to those in ResStock. We only use ResStock data from climate zone 4A, the region in which the synthetic homes are collected in our example, and we only compare the four pre-selected variables—Wall R-Value, Roof R-Value, Cooling Coefficient of Performance (COP), Heating COP—that are not static among all synthetic homes. The number of synthetic homes is $258$ while the number of climate zone 4A ResStock samples is $2000$.
@@
The results above and the tests of the generative components of our framework demonstrate a level of confidence above previous work such as Zhang et al. (2025a) and Xiao and Xu (2024). These works and others either accept the output of the LLMsor use procedural evaluation techniques.
Occlusion offers stronger empirical support that the image processing component of our pipeline—which is LLM-based—performs as expected. We additionally demonstrated statistically that properties of our generated homes are realistically distributed according to ResStock. Our results directly address and mitigate concerns about hallucination and unreliability raised in prior work (Liu et al., 2025). It is fair to be skeptical of LLMsand other MLmodels—but with our results, we clearly demonstrate our pipeline is not as susceptible to some of the issues Liu et al. (2025) raise.